Miniconference on

SMOOTH FIBRE BUNDLES AND HIGHER TORSION

Tentative Schedule

September 5–7, 2003

Friday, September 5. 15:00–16:00 X. Ma (Palaiseau). Torsion for manifolds with boundary 17:00–18:00 K. Köhler (Düsseldorf). Quaternionic torsion

Saturday, September 6. 9:00–10:00 H. Reich (Münster). Algebraic K-theory of group rings and topological cyclic homology 11:00–12:00 J. Heitsch (Chicago). Riemann Roch Grothendieck and torsion for foliations

Lunch

14:30–15:30 W. Dorabiała (Penn State). Product and sum theorems for parametrized Franz-Reidemeister torsion 16:00–17:00 M. Weiss (Aberdeen). Cohomology of the stable mapping class group 17:30–18:30 B. Williams (Notre Dame). Lefschetz-Riemann-Roch for K-theory of spaces

Sunday, September 7. 9:00–10:00 J. Klein (Wayne). Homotopical intersection theory 11:00–12:00 K. Igusa (Brandeis). Characterization of higher torsions